OpenStudy (anonymous):

How many ways can you form committee with 4 men and 11 women, if there are to be at least 4 men and twice as many women as men?

ganeshie8 (ganeshie8):

There are only 4 men available and you need atleast 4 men so you need to take all men

ganeshie8 (ganeshie8):

Since women must be twice as many as men, you need to choose exactly 8 women

ganeshie8 (ganeshie8):

THe problem simplifies to finding number of ways of choosing 4 men and 8 women

ganeshie8 (ganeshie8):

\[\large \rm \binom{4}{4} \times \binom{11}{8}\]

OpenStudy (anonymous):

Do you also have to add on the number of ways for having a committee with 8,9,10 and 11 women?

ganeshie8 (ganeshie8):

Depends on how you interpret the problem

OpenStudy (anonymous):

Like you could have a committee of 4 men, 8 or 9 or 10 or 11 women, And another set of committees with 5 men and 10 or 11 women?

ganeshie8 (ganeshie8):

Does " twice as many women as men" imply "atleast twice" ?

OpenStudy (anonymous):

Yes at least twice as many women as men.

ganeshie8 (ganeshie8):

Ohkay, then just add up other combinations too

ganeshie8 (ganeshie8):

5 men is impossible because there are only 4 men available

ganeshie8 (ganeshie8):

\[\rm \binom{4}{4} \times \left[\binom{11}{8}+ \binom{11}{9}+ \binom{11}{10}+ \binom{11}{11}\right]\] something like this ?

OpenStudy (anonymous):

Yeah but I think since the order doesn't matter wouldn't I just want to add the men to women?

ganeshie8 (ganeshie8):

Here multiplication is not about order, you need to have 4 men for sure in the committee. So you don't have any choice... Maybe assume that they are already there in the committee and select women

ganeshie8 (ganeshie8):

If you want to think of it in terms of order, try this : first find # of ways to choose four men. after that, for each of this choice, find the number of ways of choosing 8+ women

OpenStudy (anonymous):

Ok. That makes sense. I understand how to do it for four men, with the different amounts of women, Would I then add on the # of ways to have 5 men with 10 and 11 women since that is the only other possible formation since there are 6 men to choose from?

ganeshie8 (ganeshie8):

have i misread the main question.... it says there are 4 men only hm

OpenStudy (anonymous):

I made a mistake writing it. There are suppose to be 6 men, and 11 women. So with the constraint you should be able to have the choice of 5 men as well.

OpenStudy (anonymous):

Sorry...

ganeshie8 (ganeshie8):

Got you :) yes you can do the same for (5men 10/11women)

ganeshie8 (ganeshie8):

6men is not possible right ?

OpenStudy (anonymous):

Correct. Thank you so much Ganeshie!

ganeshie8 (ganeshie8):

\[\rm \binom{6}{4} \times \left[\binom{11}{8}+ \binom{11}{9}+ \binom{11}{10}+ \binom{11}{11}\right] \\~\\+ \\~\\\rm \binom{6}{5} \times \left[ \binom{11}{10}+ \binom{11}{11}\right]\]